### v1

@@ Abstract
Virtual Try-On (trying clothes virtually) is a promising application of the Generative Adversarial Network (GAN). However, it is an arduous task to transfer the desired clothing item onto the corresponding regions of a human body because of varying body size, pose, and occlusions like hair and overlapped clothes. In this paper, we try to produce photo-realistic translated images through semantic segmentation and a generative adversarial architecture-based image translation network. We present a novel image-based Virtual Try-On application VTON-IT that takes an RGB image, segments desired body part, and overlays target cloth over the segmented body region. Most state-of-the-art GAN-based Virtual Try-On applications produce unaligned pixelated synthesis images on real-life test images. However, our approach generates high-resolution natural images with detailed textures on such variant images.  [^1] Details of the implementation, algorithms and codes, are publicly available on Github: https://github.com/shuntos/VITON-IT
@@
Virtual Try On, Human Part Segmentation, Image Translation, Semantic Segmentation, Generative Adversarial Network
@@ 1 Introduction
Research and Development on Virtual Try-On applications is getting popular as the fashion e-commerce market is rapidly growing. With a virtual try-on application, customers can try the desired cloth virtually before purchasing and sellers can benefit from an increased online marketplace. In addition, this application can reduce the uncertainty of size and appearance that most online shoppers are afraid of. Problems of traditional 3D-based virtual try-on are computational complexity, tedious hardware dependent data acquisition process and less user-friendly [1]. Image based 2D virtual try-on application, if integrated into existing e-commerce or digital marketplace, will be more scalable and memory sufficient as compared to 3D approach.
With the recent development in the GAN, Image-to-image translation on the conditional setting is possible [2]. Improved discriminator and generation architecture made cross-domain high-resolution image translation possible [3]. Which makes transformation of styles and textures from one domain to another possible.
In this paper, we mainly focus on solving the problems of 2D virtual try on applications by leveraging state-of-the-art deep learning network for semantic segmentation and robust image translation network for translating the input image into target domain. In previous works like VVT [4] there are problems with semantic segmentation because of plain backgrounds in training datasets. To solve this issue we trained a UNet-like semantic segmentation architecture on diverse images manually picked from the FGV6 dataset [5]. For image translation tasks residual mapping generator and the multi-scale discriminator is used which takes a semantic mask from the segmentation network and translates it into a wrapped RGB cloth with fine details. The previous methods only worked on an image with a single person on it [6]. Therefore, to make it applicable for multi-person cases, we used a pre-trained human detection Yolov5 model [7] trained on COCO dataset [8] to generate a bounding box for each human body and crop the overlaying cloth on it. VTON-IT architecture offers pose, background, and occlusion invariant application for the online fashion industry with a wide range of applications. Rigorous testing and experiments have shown that our approach generates more visually promising overlayed images as compared to existing methods.
@@ 2 Related Works
There are a number of image based virtual try-on approaches that have been tried in the prior researches, those relevant in our study are discussed here.
@@ 2.1 VITON
Han et al. presented an image-based Virtual Try-On Network (VITON) [9]: a coarse-to-fine framework that seamlessly transferred a target clothing item in a product image to the corresponding region of a clothed person in a 2D image. The warped target clothing-to match the pose of the clothed person was generated using a thin-plate spline (TPS) transformation which is ultimately fused with the person’s image.
@@ 2.2 CP-VTON
A similar structure to VITON is adopted by CPVTON [10], but it uses a neural network to learn the spatial transformation parameters of the TPS transformation present in its Geometric Matching Module (GMM). Therefore, the GMM will create a warped cloth image and a try-on module will fuse the warped cloth image to the target image of the person, preserving the precise features of the clothes.
@@ 2.4 VTNFP
VTNFP [12] follows a three-stage design strategy by first generating warped clothing, followed by generating a body segmentation map of the person wearing the target clothing, and ending with a try-on synthesis module to fuse together all information for a final image synthesis. This method preserves target cloth and human body parts along with clothes that need no replacement.
@@ 2.5 Virtual Try-on auxiliary human segmentation
Virtual Try-On using auxiliary human segmentation [13] was based on additional improvements on existing CP-VTON. Using the human semantic segmentation prediction as an auxiliary task helped improve the virtual try-on performance. A branched architecture was proposed to simultaneously predict the try-on result and the expected segmentation mask of the generated try-on output where the target model was now wearing the in-shop cloth.
@@ 3 Proposed Approach
The network architecture proposed in this study is composed of three key components. The initial stage of the proposed architecture involves the implementation of a human body parsing network, which is followed by a body region segmentation network. Finally, an image translation network is employed to enable the wrapping of input clothing over the target body.
Figure 1: Proposed VITON-IT overview. First the human body is detected and cropped. Then, the desired body region is segmented through $U^{2}$-Net architecture and the segmented mask is fed to the image translation network to generate wrapped cloth. Finally, wrapped cloth is overlayed over the input image.
@@ 3.1 Human Parsing Network
To parse the body part for translation, the input image is given as input to Yolov5 pre-trained model trained on the Microsoft COCO dataset [8] for object detection tasks. This model outperforms existing object detection models in terms of latency and memory consumption. Model gives bounding boxes of humans in image and probability score. This allows the user to perform image translation on multiple objects. Pytorch implementation of the Yolov5-large model is used for inference. Input image of size $(640\times 640)$ is fed into the network with confidence threshold $=0.25$, Non-Maximum IOU threshold $=0.45$, and $max\_detection=10$ is used.
@@ 3.2 Human body Segmentation
$U^{2}$-Net architecture has been implemented to generate masks of desired body parts for image translation tasks. While existing backbones, such as Alexnet [14], VGG [15], ResNet [16], and DenseNet [17], are utilized for semantic segmentation tasks, their feature maps have a lower resolution. For example, ResNet reduces the size of feature maps to one-fourth of the input size. However, feature map resolution is very crucial for salient object detection (SOD). The objective of SOD is to segment the most visually appealing object in an image. Similarly, another problem of such backbones is their complex architecture because of additional feature extractor modules for extracting multi-level saliency features.
Nested UNet architecture, $U^{2}$-Net aims to go deeper while preserving feature map resolution. It can be trained from scratch and preserve the resolution of feature maps with the help of residual U-block (RSU). Top-level UNet-like architecture is followed by RSU which can extract Intra-stage multi-scale features.
Architecture: For salient object detection tasks, both local and global contextual information plays an important role. Existing feature extractor backbones have small receptive fields as they use small convolution filters of size $(1\times 1)$ or $(3\times 3)$ for the sake of computational cost and storage space. Such small receptive fields are not capable of capturing global information. Large receptive fields can capture global information from high-resolution feature maps. Dilated convolution operation results in an enlarged receptive field. However, performing multiple dilated convolution operations earlier with the original resolution needs high computational and memory resources. To tackle this problem RSU is used which has mainly three components.
@@
    Input feature map $F1(x)$ is given as input to UNet like architecture of height $L$, higher value of $L$ means deeper the network with large number of pooling layers and ranges of receptive fields and numbers of local and global features. During downsampling multi-scale features are extracted and higher resolution features maps are encoded through progressive upsampling, concatenation and convolution.
@@
    Fusion of local and global features is done through RSU. $H(x)=(F(x)+U(F(x))$ Where $F(x)=$ Intermediate feature map, $U(F(x))=$ Multi-Scale contextual information and $H(x)=$ Desired mapping of input features.
$U^{2}$-Net consists of two-level nested U structures. The first level consists of 11 well-configured RSU stages capable of extracting intra-stage multi-scale and inter-stage multi-level features. Architecture has main 3 parts: a six-stage encoder, a five-stage decoder, and a saliency map fusion module with a decoder stage and the last encoder stage. In the first four encoder stages RSU is used but in the 5th and 6th stage resolution of the feature map is relatively low, so further downsampling might lead to loss of contextual information. Thus in the 5th and 6th encoder stages dilated RSU (replaced upsampling and pooling with dilated convolution) is used to preserve the resolution of feature maps. Feature map resolution in the 4th to 6th stages is the same. In decoder stages, dilated RSU is used when each stage takes concatenated upsampled feature maps from the previous stage. Saliency map fuser inputs saliency probability maps from five decoders and the last stage (6th) encoder where each map is generated through $(3\times 3)$ convolution and sigmoid function. These probability maps are concatenated and passed through $(1\times 1)$ convolution and sigmoid function to generate the final saliency probability map $S_{fused}$.
@@ 3.3 Image Translation
Architecture: This GAN framework consists of generator $G$ and discriminator $D$ for image-image translation tasks. The task of generator $G$ is to generate an image(RGB) of cloth given a binary semantic map generated by the human body segmentation network. Whereas, discriminator $D$ tries to classify whether the generated image is real or synthesized. Dataset consists of a pair of images ($S_{i}$ and $X_{i}$) where $S_{i}$ is a mask and $X_{i}$ is the corresponding real image. This architecture works in supervised configuration to model the conditional distribution of real images with given binary masks with a min-max game. Where generator $G$ and discriminator $D$ try to win against each other. From [18], we have,
@@
This architecture uses UNet as a generator and a patch-based network as a discriminator. The generator takes 3 channel mask whereas a concatenated channel-wise semantic label map and the corresponding image are fed to Discriminator. The main component of this architecture is a coarse-to-fine generator, multi-scale discriminator, and optimized adversarial objective function. The generator consists of a global generator network $G1$ and a local enhancer network. Local generator outputs an image of resolution $4$ times larger than the original, or in other words ( 2 times of width and height) greater than the previous one. To increase the resolution of the synthesis image additional local generators can be added. For instance, the output resolution of $\{G1,G2)$ is $1024\times 2048$ whereas the output resolution of $\{G1,G2,G3\}$ is $2048\times 4096$. First global generator $G1$ is trained and then the local generator $G2$ is trained and so on, in the order of their resolution. Global generator $G1$ is trained on low-resolution images then another residual network $G2$ is added to $G1$ and a joint network is trained on higher-resolution images. Element wise sum of the feature map of $G2$ and the last feature map of $G1$ is fed into the next $G2$.
@@
Where $D_{k}$, $D_{k}^{(i)}$ are feature extractor layers of discriminator, $T$ represent total number of layers and $N$ represent number of elements per layer.
@@
This loss function works well for translating mask to cloth image with higher resolution and detail texture.
@@ 4.1 Training Human Body Segmentation network
For the training dataset, $6000$ good quality images are selected manually from FGVC6 [5] dataset and labeled using Labelme tool [19] to generate desired body masks. The average resolution of training images is $(630\times 1554)$. The model was trained through transfer learning using a pre-trained model trained on COCO dataset for general human body segmentation task with input Image size $320\times 320$ with random flip and crop. Pytorch library is used for training and inference. Adam optimizer [20] is used to train our network and its hyperparameters are set to default (initial learning rate $lr=1e-3$, $betas=(0.9,0.999)$, $eps=1e-8$, $weight\_decay=0$). And initially, the loss is set to $1$. The total number of iterations was $400000$ with a training loss of $0.109575$.
@@ 4.2 Training Image Translation Network
Dataset was prepared manually by creating a pair of real images and corresponding mask. Images are labeled using the Labelme tool to generate a semantic label and apply a series of geometrical augmentation algorithms. We prepared training pair images through data augmentation. As deep neural networks involve millions of parameters, incorporating more training data that is relevant to the domain can effectively mitigate the problem of overfitting. According to Zhao et. al [21], GAN performance is improved more by augmentations that cause spatial changes than by augmentations that just cause visual changes. Therefore, we have used several geometric augmentation techniques to augment the cloth image before training. These are namely Perspective Transform, Piecewise Affine Transform, Elastic Transformation, Shearing, and Scaling. We augmented both the image and its corresponding mask using the Imgaug library [22]. The image translation model was trained on a 3-channel input image of size $(512\times 512)$ with a batch size of $4$, without input label and instance map. The generator generates output images with the same number of image channels and shapes. Final loss consists of three components: GAN loss, discriminator-based feature matching loss, and VGG perceptual loss. After training up to $100$ epoch we got a GAN loss of $0.83$, a GAN Feature loss of $2.123$, and a VGG perceptual loss of $1.81$.
@@ 4.3 Training Setup
Both human segmentation and image translation network was trained on a Ubuntu 16.04 with 2 Nvidia GPU: GA102 [GeForce RTX 3090], 32 GB RAM and CPU: 20 core Intel(R) Core(TM) i9-10900K CPU @ 3.70GHz.
@@ 5.1 Qualitative Results
For evaluating the performance of VTON-IT through visual observation, we compared the final overlayed images with the output of CP-VTON+ [11]. Figure 4 shows that proposed virtual try-on application produces more realistic and convincing results in terms of texture transfer quality and pose preservation. Most of the existing virtual try on produce low-resolution output image. CP-VTON+ generates an output image with a fixed shape $(192\times 256)$ but our proposed approach works on high-resolution images. Through a high-resolution image translation network wrapped cloth of shape $(512\times 512)$ is generated. While experimenting with high-resolution input of shape $(2448\times 3264)$ we got a perfectly aligned natural-looking overlayed image with the same shape as input.

### v2

@@ Abstract
Virtual Try-On (trying clothes virtually) is a promising application of the Generative Adversarial Network (GAN). However, it is arduous to transfer the desired clothing item onto the corresponding regions of a human body because of varying body size, pose, and occlusions like hair and overlapped clothes. This paper aims to produce photo-realistic translated images through semantic segmentation and a generative adversarial architecture-based image translation network. We present a novel image-based Virtual Try-On application VTON-IT that takes an RGB image, segments desired body part, and overlays target cloth over the segmented body region. Most state-of-the-art GAN-based Virtual Try-On applications produce unaligned pixelated synthesis images on real-life test images. However, our approach generates high-resolution natural images with detailed textures on such variant images.  [^1] Details of the implementation, algorithms, and codes, are publicly available on Github: https://github.com/shuntos/VITON-IT
@@
Virtual Try-On, Human Part Segmentation, Image Translation, Semantic Segmentation, Generative Adversarial Network
@@ 1 Introduction
Research and Development on Virtual Try-On applications is getting popular as the fashion e-commerce market is rapidly growing. With a virtual try-on application, customers can try the desired cloth virtually before purchasing and sellers can benefit from an increased online marketplace. In addition, this application can reduce the uncertainty of size and appearance that most online shoppers are afraid of. Problems of traditional 3D-based virtual try-on are computational complexity, tedious hardware-dependent data acquisition process, and less user-friendly [1]. Image-based 2D virtual try-on applications, if integrated into existing e-commerce or digital marketplace, will be more scalable and memory efficient compared to the 3D approach.
In the recent advancements in GANs, image-to-image translation in the conditional setting has become possible [2]. Improved discriminator and generation architectures have enabled cross-domain high-resolution image translation [3], allowing for the transformation of styles and textures from one domain to another.
In this paper, we primarily address the challenges of 2D virtual try-on applications by leveraging state-of-the-art deep learning networks for semantic segmentation and robust image translation networks for translating input images into the target domain. Previous works like VVT [4] encountered issues with semantic segmentation due to plain backgrounds in training datasets. To address this, we trained a UNet-like semantic segmentation architecture on diverse images manually selected from the FGV6 dataset [5]. For image translation tasks, a residual mapping generator and multi-scale discriminator are employed, taking a semantic mask from the segmentation network and translating it into a wrapped RGB cloth with fine details. Previous methods only worked on images with a single person [6]. Thus, to make it applicable for multi-person cases, we utilized a pre-trained human detection Yolov5 model [7] trained on the COCO dataset [8] to generate bounding boxes for each human body and crop the overlaying cloth accordingly.
The VTON-IT architecture offers pose, background, and occlusion invariant applications for the online fashion industry with a wide range of applications. Rigorous testing and experiments have demonstrated that our approach generates more visually promising overlayed images compared to existing methods.
@@ 2 Related Works
Several image-based virtual try-on approaches have been tried in the prior research, those relevant in our study are discussed here.
@@ 2.1 VITON
Han et al. presented an image-based Virtual Try-On Network (VITON) [9]: a coarse-to-fine framework that seamlessly transferred a target clothing item in a product image to the corresponding region of a clothed person in a 2D image. The warped target clothing to match the pose of the clothed person was generated using a thin-plate spline (TPS) transformation which is ultimately fused with the person’s image.
@@ 2.2 CP-VTON
CPVTON [10] adopts a structure similar to VITON, but it utilizes a neural network to learn the spatial transformation parameters of the TPS transformation within its Geometric Matching Module (GMM). Consequently, the GMM generates a warped cloth image, and a try-on module fuses the warped cloth image with the target image of the person, preserving the precise features of the clothes.
@@ 2.4 VTNFP
VTNFP [12] adopts a three-stage design strategy. Initially, it generates warped clothing, followed by generating a body segmentation map of the person wearing the target clothing. Finally, it employs a try-on synthesis module to fuse all information for the final image synthesis. This method effectively preserves both the target cloth and human body parts, ensuring that clothes requiring no replacement remain intact.
@@ 2.5 Virtual Try-on auxiliary human segmentation
Virtual Try-On using auxiliary human segmentation [13] builds upon the existing CP-VTON framework by incorporating additional enhancements. Leveraging human semantic segmentation predictions as an auxiliary task significantly enhances virtual try-on performance. The proposed architecture introduces a branched design to concurrently predict the try-on outcome and the expected segmentation mask of the generated try-on output, with the target model now adorned in the in-shop cloth.
@@ 3 Proposed Approach
The network architecture proposed in this study comprises three key components. The first stage involves implementing a human body parsing network, followed by a body region segmentation network. Finally, an image translation network is employed to facilitate the wrapping of input clothing over the target body.
Figure 1: Proposed VITON-IT overview. First, the human body is detected and cropped. Then, the desired body region is segmented through $U^{2}$-Net architecture and the segmented mask is fed to the image translation network to generate wrapped cloth. Finally, the wrapped cloth is overlayed over the input image.
@@ 3.1 Human Parsing Network
To parse the body part for translation, the input image is given as input to the Yolov5 pre-trained model trained on the Microsoft COCO dataset [8] for object detection tasks. This model outperforms existing object detection models in terms of latency and memory consumption. The model gives bounding boxes of humans in image and probability score. This allows the user to perform image translation on multiple objects. Pytorch implementation of the Yolov5-large model is used for inference. Input image of size $(640\times 640)$ is fed into the network with confidence threshold $=0.25$, Non-Maximum IOU threshold $=0.45$, and $max\_detection=10$ is used.
@@ 3.2 Human body Segmentation
The $U^{2}$-Net architecture has been implemented to generate masks of desired body parts for image translation tasks. While existing backbones like AlexNet [14], VGG [15], ResNet [16], and DenseNet [17] are utilized for semantic segmentation tasks, their feature maps  [18] have lower resolutions. For instance, ResNet reduces the size of feature maps to one-fourth of the input size. However, feature map resolution is crucial for salient object detection (SOD), where the objective is to segment the most visually appealing object in an image. Additionally, these backbones often have complex architectures due to the inclusion of additional feature extractor modules for extracting multi-level saliency features.
The Nested UNet architecture, $U^{2}$-Net, aims to delve deeper while preserving feature map resolution. It can be trained from scratch and maintains the resolution of feature maps with the help of residual U-blocks (RSU). A top-level UNet-like architecture is followed by RSU, which can extract intra-stage multi-scale features.
Architecture: In salient object detection tasks, both local and global contextual information are crucial. Existing feature extractor backbones tend to have small receptive fields, using small convolution filters of size $(1\times 1)$ or $(3\times 3)$ for computational efficiency and storage space considerations. However, such small receptive fields may struggle to capture global information. Enlarging the receptive field can be achieved through dilated convolution operations. Nonetheless, multiple dilated convolution operations at the original resolution require significant computational and memory resources. To address this issue, RSU is employed, consisting of three main components.
@@
    Input feature map $F1(x)$ is given as input to UNet like the architecture of height $L$, a higher value of $L$ means deeper the network with a large number of pooling layers and ranges of receptive fields and numbers of local and global features. During downsampling multi-scale features are extracted and higher-resolution feature maps are encoded through progressive upsampling, concatenation, and convolution.
@@
    Fusion of local and global features is done through RSU. $H(x)=(F(x)+U(F(x))$ Where $F(x)=$ Intermediate feature map, $U(F(x))=$ Multi-Scale contextual information, and $H(x)=$ Desired mapping of input features.
$U^{2}$-Net consists of two-level nested U structures. The first level consists of 11 well-configured RSU stages capable of extracting intra-stage multi-scale and inter-stage multi-level features. Architecture has main 3 parts: a six-stage encoder, a five-stage decoder, and a saliency map fusion module with a decoder stage and the last encoder stage. In the first four encoder stages, RSU is used but in the 5th and 6th stage resolution of the feature map is relatively low, so further downsampling might lead to loss of contextual information. Thus in the 5th and 6th encoder stages dilated RSU (replaced upsampling and pooling with dilated convolution) is used to preserve the resolution of feature maps. Feature map resolution in the 4th to 6th stages is the same. In decoder stages, dilated RSU is used when each stage takes concatenated upsampled feature maps from the previous stage. The saliency map fuser inputs saliency probability maps from five decoders and the last stage (6th) encoder where each map is generated through $(3\times 3)$ convolution and sigmoid function. These probability maps are concatenated and passed through $(1\times 1)$ convolution and sigmoid function to generate the final saliency probability map $S_{fused}$.
@@ 3.3 Image Translation
Architecture: This GAN framework consists of generator $G$ and discriminator $D$ for image-image translation tasks. The task of generator $G$ is to generate an image(RGB) of cloth given a binary semantic map generated by the human body segmentation network. Whereas, discriminator $D$ tries to classify whether the generated image is real or synthesized. The dataset consists of a pair of images ($S_{i}$ and $X_{i}$) where $S_{i}$ is a mask and $X_{i}$ is the corresponding real image. This architecture works in supervised configuration to model the conditional distribution of real images with given binary masks with a min-max game. Where generator $G$ and discriminator $D$ try to win against each other. From [19], we have,
@@
This architecture uses UNet as a generator and a patch-based network as a discriminator. The generator takes 3 channel mask whereas a concatenated channel-wise semantic label map and the corresponding image are fed to Discriminator. The main components of this architecture are a coarse-to-fine generator, a multi-scale discriminator, and an optimized adversarial objective function. The generator consists of a global generator network $G1$ and a local enhancer network. The local generator outputs an image of resolution $4$ times larger than the original, or in other words ( 2 times of width and height) greater than the previous one. To increase the resolution of the synthesis image additional local generators can be added. For instance, the output resolution of $\{G1,G2)$ is $1024\times 2048$ whereas the output resolution of $\{G1,G2,G3\}$ is $2048\times 4096$. First global generator $G1$ is trained and then the local generator $G2$ is trained and so on, in the order of their resolution. Global generator $G1$ is trained on low-resolution images then another residual network $G2$ is added to $G1$ and a joint network is trained on higher-resolution images. Element wise sum of the feature map of $G2$ and the last feature map of $G1$ is fed into the next $G2$.
@@
Where $D_{k}$, $D_{k}^{(i)}$ are feature extractor layers of the discriminator, $T$ represents the total number of layers and $N$ represents the number of elements per layer.
@@
This loss function works well for translating mask to cloth image with higher resolution and detailed texture.
@@ 4.1 Training Human Body Segmentation network
For the training dataset, $6000$ good quality images are selected manually from FGVC6 [5] dataset and labeled using Labelme tool [20] to generate desired body masks. The average resolution of training images is $(630\times 1554)$. The model was trained through transfer learning using a pre-trained model trained on the COCO dataset for general human body segmentation task with input Image size $320\times 320$ with random flip and crop. Pytorch library is used for training and inference. Adam optimizer [21] is used to train our network and its hyperparameters are set to default (initial learning rate $lr=1e-3$, $betas=(0.9,0.999)$, $eps=1e-8$, $weight\_decay=0$). Initially, the loss is set to $1$. The total number of iterations was $400000$ with a training loss of $0.109575$.
@@ 4.2 Training Image Translation Network
The dataset was prepared manually by creating a pair of real images and a corresponding mask. Images are labeled using the Labelme tool to generate a semantic label and apply a series of geometrical augmentation algorithms. We prepared training pair images through data augmentation. As deep neural networks involve millions of parameters, incorporating more training data that is relevant to the domain can effectively mitigate the problem of overfitting. According to Zhao et. al [22], GAN performance is improved more by augmentations that cause spatial changes than by augmentations that just cause visual changes. Therefore, we have used several geometric augmentation techniques to augment the cloth image before training. These are namely Perspective Transform, Piecewise Affine Transform, Elastic Transformation, Shearing, and Scaling. We augmented both the image and its corresponding mask using the Imgaug library [23]. The image translation model was trained on a 3-channel input image of size $(512\times 512)$ with a batch size of $4$, without input label and instance map. The generator generates output images with the same number of image channels and shapes. Final loss consists of three components: GAN loss, discriminator-based feature matching loss, and VGG perceptual loss. After training up to $100$ epoch we got a GAN loss of $0.83$, a GAN Feature loss of $2.123$, and a VGG perceptual loss of $1.81$.
@@ 4.3 Training Setup
Both human segmentation and image translation network were trained on a Ubuntu 16.04 with 2 Nvidia GPU: GA102 [GeForce RTX 3090], 32 GB RAM, and CPU: 20 core Intel(R) Core(TM) i9-10900K CPU @ 3.70GHz.
@@ 5.1 Qualitative Results
For evaluating the performance of VTON-IT through visual observation, we compared the final overlayed images with the output of CP-VTON+ [11]. Figure 4 shows that the proposed virtual try-on application produces more realistic and convincing results in terms of texture transfer quality and pose preservation. Most of the existing virtual try on produce low-resolution output images. CP-VTON+ generates an output image with a fixed shape $(192\times 256)$ but our proposed approach works on high-resolution images. Through a high-resolution image translation network wrapped cloth of shape $(512\times 512)$ is generated. While experimenting with high-resolution input of shape $(2448\times 3264)$ we got a perfectly aligned natural-looking overlayed image with the same shape as input.